# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v1

@@ 3.2.2 Editing Based on Image Prompt
$\bullet$ Multi-Stage Progressive Refinement. The multi-stage approach, through phased processing, enables the model to preserve important details of the original image while enhancing the editing effects in the target regions, thereby better adapting to complex editing tasks. Point’n Move [147] additionally introduces 2D hint points as conditions and designs a two-stage self-prompting segmentation algorithm for mask refinement and merging, enabling real-time editing without requiring training for each edit. Jakub et al. [148] design a three-stage method: it begins with a set of input images and camera poses, utilizing a neural signed distance field (SDF) to reconstruct the scene surface, which guides the training of Gaussian splatting components to ensure alignment with the scene geometry. Finally, visual and geometric information is encoded into a lightweight triangle soup proxy. Edits are propagated through this intermediate structure to the mesh extracted from neural surface, thereby updating restored appearance. GaussianVTON [149] pioneers a 3D Virtual Try-On pipeline by integrating Gaussian Splatting with 2D VTON, introducing a three-stage refinement strategy to enhance consistency.
@@ 3.2.3 Style Transfer
$\bullet$ Efforts towards Objective Function. SGSST [150] designs a Simultaneously Optimized Scales loss, enabling style transfer across all scales for consistent multi-scale stylization. ReGS [151] improves stylization consistency via a Stylized Pseudo View Supervision loss for uniform appearance across views, and a Template Correspondence Matching loss to propagate style to occluded regions. WaSt-3D [152] performs style transfer directly on 3D Gaussians, using an entropy-regularized Wasserstein-2 distance to smoothly map the style scene distribution to the content scene via gradient flow, and decomposing stylization into smaller subproblems for efficiency. Multi-StyleGS [153] introduces a style loss with bipartite matching between multiple style image regions and GS points to enable local style transfer.
$\bullet$ Efforts towards Integrating Diffusion Model. InstantStyleGaussian [154] accelerates style editing by using a diffusion model to generate target style images, incorporating them into the training set, and iteratively optimizing the Gaussian splatting scene. ArtNVG [155] enhances control over content and style via the CSGO model and Tile ControlNet, and introduces an Attention-based Neighboring-View Alignment mechanism to maintain consistent colors and textures across neighboring views. Morpheus [156] employs an RGBD diffusion model with depth-guided cross-attention, feature injection, and a Warp ControlNet conditioned on composite frames to guide 3D stylization. FantasyStyle [157] proposes Controllable Stylized Distillation, achieving 3D style transfer solely through diffusion model distillation.
$\bullet$ Efforts towards VGG Feature Optimization. Saroha et al. [158] employ a pre-trained VGG model with AdaIN to transfer styles to specific views, training a 3D color module to predict new colors for each Gaussian. Similarly, StyleSplat [159] refines style transfer by applying a nearest neighbor feature matching loss between VGG features of the rendered image and the reference style image, and uses a 2D mask for selective stylization of specified objects. In contrast, StyleGaussian [160] embeds 2D VGG scene features into transformed 3D Gaussian features, combines them with reference image features, and decodes them into stylized RGB images. SemanticSplatStylization [161] integrates semantic understanding and uses VGG features to compute style loss, ensuring precise, context-aware stylization.
@@ 3.3.1 Object-level Generation
$\bullet$ Efforts towards Improving SDS. Although SDS provides guidance for 3D generation, it often suffers from low fidelity and over-smoothing. Recent works address these issues by enhancing geometric accuracy, detail preservation, and stability. (a) Text-to-3D. StableDreamer [181] formulates the SDS prior and L2 reconstruction loss as equivalent, reducing multi-face artifacts and improving geometry. Li et al. [182] propose Guided Consistency Sampling integrated with 3DGS to enhance detail and fidelity. DreamMapping [183] introduces Variational Distribution Mapping, accelerating distribution modeling by treating rendered images as degraded diffusion outputs. HumanGaussian [184] designs a structure-aware SDS for joint optimization of human appearance and geometry. LucidDreamer [41] mitigates over-smoothing via Interval Score Matching (ISM) with deterministic diffusion trajectories, while TSM [185] extends ISM by generating two reverse DDIM [186] paths to reduce cumulative errors and pseudo-ground truth inconsistencies. GaussianMotion [187] employs adaptive score distillation to balance realism and smoothness. (b) Image-to-3D. Basak et al. [188] introduce a frequency-based distillation loss, extracting low-frequency geometry from 3D diffusion priors and refining high-frequency texture with 2D diffusion. GECO [189] adopts a two-stage pipeline combining Variational Score Distillation [190] and multi-view consistency refinement, enabling sub-second high-quality generation. DreamPhysics [191] leverages video diffusion priors to learn material properties and applies motion distillation sampling to emphasize dynamic cues.
@@
B. Feed-Forward-based Methods: Although these optimization-based methods have made significant progress in 3D content generation, they still face challenges such as lengthy optimization times. To address these issues, researchers have begun to shift towards feed-forward methods [197][198], which involve training on large-scale datasets to directly generate 3D assets.
@@ 4 Performance Comparison
TABLE II: Quantitative 3D instance segmentation on Replica[279], and LERF-Mask[22]. Following [20], we report mIoU and F-score on Replica, and mIoU and mBIoU on LERF-Mask.
@@
TABLE III: Quantitative 3D interactive segmentation on NVOS [281], SPIn-NeRF[66] in terms of mIoU and mAcc.
@@
Method Venue Train Input mIoU(%) mAcc(%) mIoU(%) mAcc(%)
@@ 4.1 Performance Benchmarking: 3DGS Segmentation
Tanks and Temples [287] [TOG’17] 14 - It includes individual objects (e.g., ‘Tank’, ‘Train’) and large indoor scenes (e.g., ‘Auditorium’, ‘Museum’).
GL3D [288] [ACCV’18] 543 230 It contains 125,623 high-resolution images, most of which were captured by drones from multiple scales and perspectives with significant geometric overlap. The images cover urban areas, rural regions, and scenic spots.
LLFF [289] [TOG’19] 32 25 Using the COLMAP structure from motion implementation to compute poses for our real images.
@@
  Methods Venue Category Condition 2D Model $\text{CLIP}_{dir}$ $\uparrow$ $\text{CLIP}_{T2I}$ $\uparrow$ $\text{CLIP}_{I2I}$ $\uparrow$ $\text{DINO}_{sim}$ $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FPS $\downarrow$
@@
Eyal et al. [132] [ArXiv’24] Multi-View Consistency Text Stable Diffusion [55] +26.00% +9.09% +10.50% +43.44% - +2.37% - -
@@ 4.2 Performance Benchmarking: 3DGS Editing
DTU [286] containing 80 scenes of large variability. Each scene consists of 49 or 64 accurate camera positions and reference structured light scans, all acquired by a 6-axis industrial robot.
@@
NYUdepth [298] [ECCV’12] Image-to-3D 464 It contains 1449 RGBD images, capturing 464 diverse indoor scenes, with detailed annotations.
@@
LOM [303] [AAAI’24] Image-to-3D 5 Includes five real-world scenes, each with 25–48 sRGB images captured in difficult lighting.
G-objaverse [304] [ECCV’24] Image&Text-to-3D 280K 10 general classes which gives about 280K samples
@@
DiffusionGS [214] [ArXiv’24] Optimization Diffusion Image FLUX & SD [55] 22.07 0.854 0.111 - 11.52 -
@@ 4.3 Performance Benchmarking: 3DGS Generation
LOM [303] comprises five real-world scenes (“buu”, “chair”, “sofa”, “bike”, and “shrub”), each containing 25 to 48 sRGB images captured by a DJI Osmo Action 3 camera under adverse lighting conditions, including low light and overexposure.
@@
Inception Score evaluate the quality of generated images using Inception-V3 [295].
@@ 5 Future Directions
$\bullet$ Faithful Evaluation Metrics. Current evaluation protocols often inherit metrics from traditional 2D or NeRF-based tasks, which may not align well with 3DGS goals. Developing reliable and interpretable metrics specifically designed for evaluating application tasks from a 3D Gaussian perspective is essential to track progress and guide future research. Existing metrics primarily focus on 2D projection image-level accuracy, lacking meaningful insights into the inherent 3D geometric and semantic consistency of generated or edited scenes. Establishing 3D-aware metrics is promising.
@@
$\bullet$ Combining 3D Foundation Model. Emerging 3D foundation models such as VGGT [317] and FLARE [318] provide robust representations by pretraining on diverse 3D tasks, including depth estimation, correspondence, and pose prediction from multi-view images. These models encode rich geometric priors that can benefit various downstream applications. Future work may incorporate them into 3DGS to improve generalization and efficiency.
@@ 6 Conclusion
This survey systematically explores recent advances in utilizing 3D Gaussian Splatting (3DGS) for downstream application tasks, representing a pioneering effort in summarizing this emerging field. We first introduce essential background knowledge, covering fundamental concepts and foundational models. Subsequently, we comprehensively review over 150 representative methods across several key tasks, including segmentation, editing, generation, and other related application tasks, providing a structured categorization from a technical perspective. Furthermore, we present a thorough analysis of existing benchmarks and conduct fair comparisons of representative models clearly. Lastly, we highlight current challenges and outline promising future directions to inspire continued research efforts toward enhanced high-level tasks.

### v2

@@ 3.2.2 Editing Based on Image Prompt
$\bullet$ Multi-Stage Progressive Refinement. The multi-stage approach, through phased processing, enables the model to preserve important details of the original image while enhancing the editing effects in the target regions, thereby better adapting to complex editing tasks. Point’n Move [153] additionally introduces 2D hint points as conditions and designs a two-stage self-prompting segmentation algorithm for mask refinement and merging, enabling real-time editing without requiring training for each edit. Szymkowiak et al. [154] design a three-stage method: it begins with a set of input images and camera poses, utilizing a neural signed distance field (SDF) to reconstruct the scene surface, which guides the training of Gaussian splatting components to ensure alignment with the scene geometry. Finally, visual and geometric information is encoded into a lightweight triangle soup proxy. Edits are propagated through this intermediate structure to the mesh extracted from neural surface, thereby updating restored appearance. GaussianVTON [155] pioneers a 3D Virtual Try-On pipeline by integrating Gaussian Splatting with 2D VTON, introducing a three-stage refinement strategy to enhance consistency.
@@ 3.2.3 Style Transfer
$\bullet$ Efforts toward Objective Function. SGSST [156] designs a Simultaneously Optimized Scales loss, enabling style transfer across all scales for consistent multi-scale stylization. ReGS [157] improves stylization consistency via a Stylized Pseudo View Supervision loss for uniform appearance across views, and a Template Correspondence Matching loss to propagate style to occluded regions. WaSt-3D [158] performs style transfer directly on 3D Gaussians, using an entropy-regularized Wasserstein-2 distance to smoothly map the style scene distribution to the content scene via gradient flow, and decomposing stylization into smaller subproblems for efficiency. Multi-StyleGS [159] introduces a style loss with bipartite matching between multiple style image regions and GS points to enable local style transfer.
$\bullet$ Efforts toward Integrating Diffusion Model. InstantStyleGaussian [160] accelerates style editing by using a diffusion model to generate target style images, incorporating them into the training set, and iteratively optimizing the Gaussian splatting scene. ArtNVG [161] enhances control over content and style via the CSGO model and Tile ControlNet, and introduces an Attention-based Neighboring-View Alignment mechanism to maintain consistent colors and textures across neighboring views. Morpheus [162] employs an RGBD diffusion model with depth-guided cross-attention, feature injection, and a Warp ControlNet conditioned on composite frames to guide 3D stylization. FantasyStyle [163] proposes Controllable Stylized Distillation, achieving 3D style transfer solely through diffusion model distillation.
$\bullet$ Efforts toward VGG Feature Optimization. Saroha et al. [164] employ a pre-trained VGG model with AdaIN to transfer styles to specific views, training a 3D color module to predict new colors for each Gaussian. Similarly, StyleSplat [165] refines style transfer by applying a nearest neighbor feature matching loss between VGG features of the rendered image and the reference style image, and uses a 2D mask for selective stylization of specified objects. In contrast, StyleGaussian [166] embeds 2D VGG scene features into transformed 3D Gaussian features, combines them with reference image features, and decodes them into stylized RGB images. SemanticSplatStylization [167] integrates semantic understanding and uses VGG features to compute style loss, ensuring precise, context-aware stylization.
@@ 3.3.1 Object-level Generation
$\bullet$ Efforts toward Improving SDS. Although SDS provides guidance for 3D generation, it often suffers from low fidelity and over-smoothing. Recent works address these issues by enhancing geometric accuracy, detail preservation, and stability. (a) Text-to-3D. StableDreamer [187] formulates the SDS prior and L2 reconstruction loss as equivalent, reducing multi-face artifacts and improving geometry. Li et al. [188] propose Guided Consistency Sampling integrated with 3DGS to enhance detail and fidelity. DreamMapping [189] introduces Variational Distribution Mapping, accelerating distribution modeling by treating rendered images as degraded diffusion outputs. HumanGaussian [190] designs a structure-aware SDS for joint optimization of human appearance and geometry. LucidDreamer [41] mitigates over-smoothing via Interval Score Matching (ISM) with deterministic diffusion trajectories, while TSM [191] extends ISM by generating two reverse DDIM [192] paths to reduce cumulative errors and pseudo-ground truth inconsistencies. GaussianMotion [193] employs adaptive score distillation to balance realism and smoothness. (b) Image-to-3D. Basak et al. [194] introduce a frequency-based distillation loss, extracting low-frequency geometry from 3D diffusion priors and refining high-frequency texture with 2D diffusion. GECO [195] adopts a two-stage pipeline combining Variational Score Distillation [196] and multi-view consistency refinement, enabling sub-second high-quality generation. DreamPhysics [197] leverages video diffusion priors to learn material properties and applies motion distillation sampling to emphasize dynamic cues.
@@
B. Feed-Forward-based Methods: Although these optimization-based methods have made significant progress in 3D content generation, they still face challenges such as lengthy optimization times. To address these issues, researchers have begun to shift toward feed-forward methods [203][204], which involve training on large-scale datasets to directly generate 3D assets.
@@ 4 Performance Comparison
TABLE II: Quantitative 3D instance segmentation on Replica[285] and LERF-Mask[22]. Following [20], we report mIoU and F-score on Replica, and mIoU and mBIoU on LERF-Mask.
@@
TABLE III: Quantitative 3D interactive segmentation on NVOS [287] and SPIn-NeRF[67] in terms of mIoU and mAcc.
@@
Method Venue Train Input Prompt mIoU(%) mAcc(%) mIoU(%) mAcc(%)
@@ 4.1 Performance Benchmarking: 3DGS Segmentation
Tanks and Temples [293] [TOG’17] 14 - It includes individual objects (e.g., “Tank”, “Train”) and large indoor scenes (e.g., “Auditorium”, “Museum”).
GL3D [294] [ACCV’18] 543 230 It contains 125,623 high-resolution images, most of which are captured by drones from multiple scales and perspectives with significant geometric overlap. The images cover urban areas, rural regions, and scenic spots.
LLFF [295] [TOG’19] 32 25 Using the COLMAP structure from motion implementation to compute poses for real images.
@@
  Methods Venue Category Condition Foundation Model $\text{CLIP}_{dir}$ $\uparrow$ $\text{CLIP}_{T2I}$ $\uparrow$ $\text{CLIP}_{I2I}$ $\uparrow$ $\text{DINO}_{sim}$ $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FPS $\downarrow$
@@
Gomel et al. [138] [ArXiv’24] Multi-View Consistency Text Stable Diffusion [55] +26.00% +9.09% +10.50% +43.44% - +2.37% - -
@@ 4.2 Performance Benchmarking: 3DGS Editing
DTU [292] contains 80 scenes of large variability. Each scene consists of 49 or 64 accurate camera positions and reference structured light scans, all acquired by a 6-axis industrial robot.
@@
NYUdepth [304] [ECCV’12] Image-to-3D 464 It contains 1449 RGB-D images, capturing 464 diverse indoor scenes with detailed annotations.
@@
[…]
G-objaverse [310] [ECCV’24] Image&Text-to-3D 280K 10 general classes which gives about 280K samples.
DL3DV-10K [311] [CVPR’24] Image-to-3D 10K Large-scale scene dataset that contains both indoor and outdoor scenarios.
@@
DiffusionGS [220] [ArXiv’24] Optimization Diffusion Image FLUX [57] & SD [55] 22.07 0.854 0.111 - 11.52 -
@@ 4.3 Performance Benchmarking: 3DGS Generation
[…]
@@
DL3DV-10K [311] comprises 51.2M frames from 10,510 videos across 65 locations, covering diverse indoor and outdoor scenes with varying reflection, transparency, and lighting conditions.
@@
Inception Score evaluates the quality of generated images using Inception-V3 [301].
@@ 5 Future Directions
$\bullet$ Faithful Evaluation Metrics. Current evaluation protocols often inherit metrics from traditional 2D or NeRF-based tasks, which may not align well with 3DGS goals. Developing reliable and interpretable metrics specifically designed for evaluating application tasks from a 3D Gaussian perspective is essential to track progress and guide future research. Existing metrics mainly evaluate accuracy after projection onto 2D images, lacking meaningful insights into the inherent 3D geometric and semantic consistency of generated or edited scenes.
@@
$\bullet$ Combining 3D Foundation Model. Emerging 3D foundation models such as VGGT [323] and FLARE [324] demonstrate robust, generalizable representations by pretraining on diverse 3D tasks, including depth estimation, point cloud reconstruction, camera prediction, and 3D point tracking from multi-view images. These models encode rich geometric priors that can benefit various downstream applications. Future work may incorporate them into 3DGS to improve generalization and efficiency.
@@ 6 Conclusion
This survey systematically explores recent advances in utilizing 3D Gaussian Splatting (3DGS) for downstream application tasks, representing a pioneering effort in summarizing this emerging field. We first introduce essential background knowledge, covering fundamental concepts and foundational models. Subsequently, we comprehensively review over 200 representative methods across several key tasks, including segmentation, editing, generation, and other related application tasks, providing a structured categorization from a technical perspective. Furthermore, we present a thorough analysis of existing benchmarks and conduct fair comparisons of representative models clearly. Lastly, we highlight current challenges and outline promising future directions to inspire continued research efforts toward enhanced high-level tasks.